A cage contains 4 black and 5 white mice that are available for a biology experiment. Three mice are selected at random. What is the probability that at least one is black and at least one is white?

Answer by ewatrrr(24785)

P(at least one white) = 1 - P(no white) = 1- (4C3/9C3)